OpenStudy (anonymous):

how to prove (sin^3x)(1+cos^2x)=sin

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

\[(\sin^3x)(1+\cos^2x)=\sin x\]\[(1+\cos^2x)={\sin x \over \sin^3x}\]\[(1+\cos^2x)={1 \over \sin^2x}\]

OpenStudy (mertsj):

You may have posted it incorrectly. I don't believe it is true. For example, substitute pi/6 for x and you will get: 1/8(1+3/4)=1/2. Clearly not true.